please tutor can you please help me figure out this problem. Thank you!
The 3rd planet from the sun, earth, is said to be 4.54 × 10^9 years old. Modern humans have been walking on this planet for 3.0 × 10^5 years**. What percentage of time has the modern human occupied the earth? Express your answer as a complete sentence using scientific notation.

Answer by Alan3354(69443)

The 3rd planet from the sun, earth, is said to be 4.54 × 10^9 years old. Modern humans have been walking on this planet for 3.0 × 10^5 years**. What percentage of time has the modern human occupied the earth?
------------------------
3*10^5 is 3e5. Commonly used and used by some computer programming.
---------------------------------------
3e5/4.54e9 = 3/4.54e4
As % = 100*3/4.54e4 = 3/4.54e2 = 3/454 %